annikalo26:

The Smiths are sending out a batch of holiday cards. Emily signed of the cards. Then Josh signed of the remaining cards. Tatiana signed 20% of what Josh had left. If there were 75 cards in all, how many cards are left to be signed?

Laylalyssa:

Are there any answer choices?

annikalo26:

No.

supie:

Ok so since Emily signed 2/5 of the cards, we have to multiply (75) by the number that was signed so 75x2/5=30 75(2/5)=30 Now, we have to subtract so, 75-30=45 so, 45 cards are left unsigned Then, we have to do the method we did first again which is ^^ but the numbers will change since the new total number is 45 so, 45x 1/9= 5 then, 45-5= 40 Now, we have to do the second step again in decimal form . . . 20% = 20/100 = .2 so, now 40x0.2=8 40-8=? Tell me if that doesn't make sense or if you need more help.

annikalo26:

I got it thanks.

supie:

Ok, no problem, you can click the Best Response button by my reply its your choice tho.

annikalo26:

K.

supie:

Thanks!

annikalo26:

Welcome.